Fifteen pupils from school B have mean IQ of 110 and a variance of 64, while 20 pupils from school C show a mean IQ of 114 and a variance of 78. What is the z-computed value?


1
Expert's answer
2022-06-06T14:37:54-0400
"z=\dfrac{\bar{X}_1-\bar{X}_2}{\sqrt{\sigma_1^2/n_1+\sigma_2^2/n_2}}"

"=\dfrac{110-114}{\sqrt{64/15+78/20}}"

"=-1.3997"